# Continuous fields of interval algebras

### v1

@@ A. The existence
Assume that $(\phi_{x})_{x}$ satisfies (ii). Let $a\in\mathcal{A}$. By assumption, we know that for any $x\in X$ and any $\epsilon>0$, there exists $\chi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ and an open neighborhood $V$ of $x$ such that $\|\chi(a)(v)-\phi_{v}(a(v))\|<\epsilon$ for all $v\in V$. Consequently, for any $x\in X$ and any $\epsilon>0$, there exist $b:=\chi(a)\in\mathcal{B}$ and an open neighborhood $V$ of $x$ such that $\phi_{v}(a(v))\in V_{\epsilon}^{b}$ for all $v\in V$. By the continuity criterion of Section 2, we deduce that $(\phi_{x}(a(x)))_{x}\in\mathcal{B}$ for any $a\in\mathcal{A}$. The equivalence between (i) and (ii) now follows by Section 2.
Assume that $(\phi_{x})_{x}$ satisfies (ii), again. Consider a basic open set $V_{(F,\epsilon)}^{\chi}$ of $\tau_{(\mathcal{A}\rightarrow\mathcal{B})}$. Let $v\in f^{-1}(V_{(F,\epsilon)}^{\chi})$. We have $\phi_{v}\simeq_{(F_{v},\epsilon)}\chi_{v}$. Write $C:=\max_{a\in F}\|\phi_{v}(a(v))-\chi(a)(v)\|$. Observe that $C<\epsilon$. Fix $\delta>0$ be such that $3\delta+C<\epsilon$. Using the assumption, we know that there exist $\psi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ and an open neighborhood $W$ of $v$ such that $\phi_{w}\simeq_{(F_{w},\delta)}\psi_{w}$ for all $w\in W$. We deduce that $\chi_{v}\simeq_{(F_{v},C+\delta)}\psi_{v}$. By Lemma 3.1, we can find an open neighborhood $Z$ of $v$ such that $Z\subseteq W\cap f^{-1}(V_{(F,\epsilon)}^{\chi})$ and such that $\chi_{z}\simeq_{(F_{z},C+2\delta)}\psi_{z}$ for all $z\in Z$. Finally, we compute that $\phi_{z}\simeq_{(F_{z},C+3\delta<\epsilon)}\chi_{z}$ for any $z\in Z$. As a result, we see that $Z\subseteq f^{-1}(V_{(F,\epsilon)}^{\chi})$ and we conclude that $(\phi_{x})_{x}$ satisfies (iii).
Assume that $(\phi_{x})_{x}$ satisfies (iii) and that $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ satisfies (PP). Let $x\in X$, let $F\subseteq\mathcal{A}$ be a finite set and let $\epsilon>0$. Using (PP), we can find some $\chi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ such that $\chi_{x}=\phi_{x}$. Consider $V:=f^{-1}(X_{(F,\epsilon)}^{\chi})$ where $f$ denotes the assignment $x\mapsto\phi_{x}$. By continuity of $f$, $V$ is an open neighborhood of $x$ and hence, $\phi_{x}\in V_{(F,\epsilon)}^{\chi}$. This shows that $(\phi_{x})_{x}$ satisfies (ii). ∎
@@
$V_{\epsilon}^{\chi}:=\{\phi_{v}\in\operatorname{Hom}_{C^{*}}(A_{v},B_{v})\mid v\in V\text{ and }d_{v}(\phi_{v},\chi_{v})<\epsilon\}$
where $V\subseteq X$ is an open set, $\epsilon>0$, $\chi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ and $d_{v}$ is any complete metric on $\operatorname{Hom}_{C^{*}}(A_{v},B_{v})$ as constructed before.
[…]
Let us now turn to metrizability of $\tau_{(\mathcal{A}\rightarrow\mathcal{B})}$. Assume that $(X,\{A_{x}\}_{x},\mathcal{A})$ is a continuous field of separable $C^{*}$-algebras over a compact Hausdorff metric space $(X,d_{X})$, whose fibers are separable $C^{*}$-algebras. The Hom-bundle $(\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x}),\tau_{(\mathcal{A}\rightarrow\mathcal{B})})$ becomes metrizable as follows.
Fix any complete metric $d_{x}$ on the set $\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$ for any $x\in X$. For any $\phi_{v},\psi_{w}\in\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$, we set
$d(\phi_{v},\psi_{w})=\inf_{\epsilon>0}\{\phi_{v},\psi_{w}\in V_{\epsilon}^{\chi}\text{ for some }V\subseteq X\text{and }\chi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})\}+d_{X}(v,w).$
It can be checked that $d$ is a well-defined metric on $\operatorname{Hom}_{C^{*}}(A,B)$.
[…]
@@
$\bullet$ Application of the selection theorem. Let $(X,d_{X})$ be any compact Hausdorff metric space of dimension at most one. Let $(X,\{A_{x}\}_{x},\mathcal{A})$ and $(X,\{B_{x}\}_{x},\mathcal{B})$ be continuous fields of unital separable $\operatorname{AI}$-algebras.
Fix metrics $\{d_{x}\}_{x}$ and $d$ on $\{\operatorname{Hom}_{C^{*}}(A_{x},B_{x})\}_{x}$ and $\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$ respectively as in  A. Recall that $d$ induces $\tau_{(\mathcal{A}\rightarrow\mathcal{B})}$ and $d_{x}$ induces $\tau_{pn}$. Also that the restriction of $d$ on $\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$ is $d_{x}$. In particular, $(\operatorname{Hom}_{C^{*}}(A_{x},B_{x}),d_{\mid})$ is a complete space.
Let $\alpha\colon\operatorname{Cu}(\mathcal{A})\rightarrow\operatorname{Cu}(\mathcal{B})$ be any $\operatorname{Cu}$-morphism which descends to the fibers and such that $\alpha_{x}([1_{A_{x}}])=[1_{B_{x}}]$. For any $x\in X$, let us consider
$\Theta_{\alpha_{x}}:=\{\phi_{x}\in\operatorname{Hom}_{C^{*}}(A_{x},B_{x})\mid\operatorname{Cu}(\phi_{x})=\alpha_{x}\}$
By Section 3, we know that $\Theta_{\alpha_{x}}$ is a non-empty closed subset of $(\operatorname{Hom}_{C^{*}}(A_{x},B_{x}),d_{\mid})$, and hence, $\Theta_{\alpha_{x}}$ is a non-empty closed subset $(\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x}),d)$. Therefore, $(\Theta_{\alpha_{x}},d_{\mid})$ is a complete space for any $x\in X$. Further, by [5, Theorem 2.4], we know that $(\Theta_{\alpha_{x}},d_{\mid})$ is path-connected for any $x\in X$. Consider the following multi-valued map
$\begin{array}[]{ll}\Theta_{\alpha}\colon X\longrightarrow\mathcal{S}_{\alpha}\\ \quad\quad x\longmapsto\Theta_{\alpha_{x}}\end{array}$
where $\mathcal{S}_{\alpha}:=\{\Theta_{\alpha_{x}}\mid x\in X\}$ is a family of subsets of $(\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x}),d)$.
The following lemma will be needed in the proof of the lower-semicontinuity of $\Theta_{\alpha}$.
@@
Let $(X,\{A_{x}\}_{x},\mathcal{A})$ and $(X,\{B_{x}\}_{x},\mathcal{B})$ be a continuous field of $C^{*}$-algebras over the same space $X$. Let $\alpha,\beta\colon\operatorname{Cu}(\mathcal{A})\rightarrow\operatorname{Cu}(\mathcal{B})$ be $\operatorname{Cu}$-morphisms descending to the fibers. Assume that there exists $x\in X$ such that $\alpha_{x}=\beta_{x}$.
Then for any finite set $G\subseteq\operatorname{Cu}(\mathcal{A})$, there exists an open neighborhood $V$ of $x$ such that $\alpha_{v}\simeq_{G_{v}}\beta_{v}$.
@@
The map $\Theta_{\alpha}\colon X\longrightarrow\mathcal{S}_{\alpha}$ is lower-semicontinuous whenever $(X,\{A_{x}\}_{x},\mathcal{A})$ is a locally trivial field and $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ satisfies $\mathrm{(PP)}$.
@@
Let $V_{\epsilon}^{\chi}$ be a basic open set of $\tau_{(\mathcal{A}\rightarrow\mathcal{B})}$. We aim to show that $\{x\in X\mid\Theta_{\alpha_{x}}\cap V_{\epsilon}^{\chi}\neq\emptyset\}$ is open in $X$. Let $x\in X$ be such that $\Theta_{\alpha_{x}}\cap V_{\epsilon}^{\chi}\neq\emptyset$. It suffices to prove that there exists an open neighborhood $Y$ of $x$ such that $Y\subseteq\{x\in X\mid\Theta_{\alpha_{x}}\cap V_{\epsilon}^{\chi}\neq\emptyset\}$. We know that $x\in V$ and that there exists $\phi_{x}\in\Theta_{\alpha_{x}}$ such that $d_{\mid}(\phi_{x},\chi_{x})<\epsilon$. Choose $\epsilon^{\prime}>0$ such that $d_{\mid}(\phi_{x},\chi_{x})<\epsilon^{\prime}<\epsilon$. By Lemma 3.1 and (PP), we can find $\psi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ together with an open neighborhood $W$ of $x$ such that $\psi_{x}=\phi_{x}$ and $d_{\mid}(\psi_{w},\chi_{w})<\epsilon^{\prime}$ for all $w\in W$.
Fix $\delta>0$ small enough such that $\delta+\epsilon^{\prime}<\epsilon$. By Section 3 together with the local triviality of $(X,\{A_{x}\}_{x},\mathcal{A})$, we can find a finite set $G\subseteq\operatorname{Cu}(\mathcal{A})$ large enough and an open neighborhood $Z$ of $x$ such that for all $z\in Z$ we have the following.
$\quad\operatorname{Cu}(\psi_{z})\simeq_{G_{z}}\alpha_{z}\Rightarrow\exists\phi_{z}\in\Theta_{\alpha_{z}}\mid\psi_{z}\simeq_{(F_{z},\delta)}\phi_{z}$
Finally, since $\operatorname{Cu}(\psi_{x})=\operatorname{Cu}(\phi_{x})=\alpha_{x}$, we can apply Lemma 3.2 to find an open neighborhood $\tilde{Z}$ of $x$ such that $\operatorname{Cu}(\psi_{z})\simeq_{G_{z}}\alpha_{z}$ for all $z\in\tilde{Z}$. Putting everything together, we deduce that for any $z\in Z\cap\tilde{Z}$, there exists $\phi_{z}\in\Theta_{\alpha_{z}}\mid\psi_{z}\simeq_{(F_{z},\delta)}\phi_{z}$. Subsequently, we compute that
$d_{\mid}(\phi_{z},\chi_{z})\leq d_{\mid}(\phi_{z},\psi_{z})+d_{\mid}(\psi_{z},\chi_{z})\leq\delta+\epsilon^{\prime}<\epsilon$
for any $z\in W\cap Z\cap\tilde{Z}$. As a conclusion, for any $y\in Y:=W\cap Z\cap\tilde{Z}\cap V$, we have $\Theta_{\alpha_{y}}\cap V_{\epsilon}^{\chi}\neq\emptyset$ which ends the proof. ∎
As a consequence, we can invoke the selection theorem to obtain the following existence result.
[…]
Let $A,B$ be a unital interval algebras
For any scaled $\operatorname{Cu}$-morphism $\alpha\colon\operatorname{Cu}(C(X)\otimes A)\rightarrow\operatorname{Cu}(C(X)\otimes B)$ descending to the fibers, there exists $\phi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ such that $\operatorname{Cu}(\phi)=\alpha$. A fortiori, $\operatorname{Cu}(\phi_{x})=\alpha_{x}$ for any $x\in X$.
Proof.
It is immediate that the triviality of $(X,\{A_{x}\}_{x},\mathcal{A})$ and $(X,\{B_{x}\}_{x},\mathcal{B})$ implies local triviality of $(X,\{A_{x}\}_{x},\mathcal{A})$ and (PP) of $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$. Moreover, the fact that $A_{x}$ are interval algebras combined with the (local) triviality of $(X,\{A_{x}\}_{x},\mathcal{A})$, we deduce the uniform path property from [5, Theorem 2.3] that we will explicitly recall and use later. (See  B). Lastly, as a consequence of  A, we are able to invoke the selection theorem yielding to a continuous selection $\phi$ for $\Theta_{\alpha}$. That $\phi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ follows from the continuity criterion of  A. That $\operatorname{Cu}(\phi)=\alpha$ follows from construction. ∎
The proof reveals that the local triviality of $(X,\{A_{x}\}_{x},\mathcal{A})$ implies the uniform path property of $\mathcal{S}_{\alpha}$. The restriction to trivial continuous fields was imposed solely to obtain (PP). Consequently, establishing (PP) for certain locally trivial fields would lead to a more general existence result.
[…]
We thank M. Dădărlat for pointing out that the Mayer-Vietoris sequence applied to $\mathrm{K}$-theory could provide an obstruction to (PP) for locally trivial fields over the circle. Let us construct a specific example.
@@
Finally, let $V,W$ be proper closed subintervals of $\mathbb{T}$ such that $V\cup W=\mathbb{T}$. It is well known that $\mathcal{A}\simeq\mathcal{A}(V)\oplus_{\mathcal{A}(V\cap W)}\mathcal{A}(W)$ and $\mathcal{B}\simeq\mathcal{B}(V)\oplus_{\mathcal{B}(V\cap W)}\mathcal{B}(W)$. (See e.g. [4] for the sheaves picture of continuous fields.)
The Mayer-Vietoris sequence yields $\mathrm{K}_{0}(\mathcal{A})\simeq\mathrm{K}_{0}(\mathcal{B})$ which leads to a contradiction. As a consequence, $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ does not have (PP).
[…]
Let $\mathcal{A},\mathcal{B}$ be continuous fields of interval algebras over the interval. Under which conditions $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ has (PP)?

### v2

@@ A. The existence
Assume that $(\phi_{x})_{x}$ satisfies (ii). Let $a\in\mathcal{A}$. By assumption, we know that for any $x\in X$ and any $\epsilon>0$, there exists $\chi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ and an open neighborhood $V$ of $x$ such that $\|\chi(a)(v)-\phi_{v}(a(v))\|<\epsilon$, for all $v\in V$. Consequently, for any $x\in X$ and any $\epsilon>0$, there exist $b:=\chi(a)\in\mathcal{B}$ and an open neighborhood $V$ of $x$ such that $\phi_{v}(a(v))\in V_{\epsilon}^{b}$ for all $v\in V$. By the continuity criterion of Section 2, we deduce that $(\phi_{x}(a(x)))_{x}\in\mathcal{B}$ for any $a\in\mathcal{A}$. The equivalence between (i) and (ii) now follows from Section 2.
Assume that $(\phi_{x})_{x}$ satisfies (ii), again. Consider a basic open set $V_{(F,\epsilon)}^{\chi}$ such that $f^{-1}(V_{(F,\epsilon)}^{\chi})\neq\emptyset$. Let $v\in f^{-1}(V_{(F,\epsilon)}^{\chi})$. We have $\phi_{v}\simeq_{(F_{v},\epsilon)}\chi_{v}$. Write $C:=\max_{a\in F}\|\phi_{v}(a(v))-\chi(a)(v)\|$. Observe that $C<\epsilon$. Fix $\delta>0$ be such that $3\delta+C<\epsilon$. Using the assumption, we know that there exist $\psi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ and an open neighborhood $W$ of $v$ such that $\phi_{w}\simeq_{(F_{w},\delta)}\psi_{w}$ for all $w\in W$. We deduce that $\chi_{v}\simeq_{(F_{v},C+\delta)}\psi_{v}$. By Lemma 3.6, we can find an open neighborhood $Z$ of $v$ such that $\chi_{z}\simeq_{(F_{z},C+2\delta)}\psi_{z}$ for all $z\in Z$. Finally, we compute that $\phi_{y}\simeq_{(F_{y},C+3\delta<\epsilon)}\chi_{y}$ for any $y\in Z\cap W$. As a result, we see that $Z\cap W\subseteq f^{-1}(V_{(F,\epsilon)}^{\chi})$ and we conclude that $(\phi_{x})_{x}$ satisfies (iii).
Assume that $(\phi_{x})_{x}$ satisfies (iii) and that $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ satisfies $\rm(PP)$. Let $x\in X$, let $F\subseteq\mathcal{A}$ be a finite set and let $\epsilon>0$. Using $\rm(PP)$, we can find some $\chi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ such that $\chi_{x}=\phi_{x}$. By the continuity of $f$, we deduce that $V:=f^{-1}(X_{(F,\epsilon)}^{\chi})$ is an open neighborhood of $x$ such that $f(v)\in V_{(F,\epsilon)}^{\chi}$ for all $v\in V$, yielding (ii). ∎
@@
$V_{\epsilon}^{\chi}:=\{(\xi,x)\in\bigsqcup_{x\in X}\operatorname{Hom}_{C^{*}}(A_{x},B_{x})\mid x\in V\text{ and }d_{x}(\xi,\chi_{x})<\epsilon\}$
where $V\subseteq X$ is an open set, $\epsilon>0$, $\chi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ and $d_{x}$ is any complete metric on $\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$ as constructed before.
3.9.
Let us now turn to metrizability of $\tau_{(\mathcal{A}\rightarrow\mathcal{B})}$. Assume that $(X,\{A_{x}\}_{x},\mathcal{A})$ is a continuous field of separable $C^{*}$-algebras over a compact Hausdorff metric space $(X,d_{X})$. The Hom-bundle $(\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x}),\tau_{(\mathcal{A}\rightarrow\mathcal{B})})$ becomes metrizable as follows.
Fix any complete metric $d_{x}$ on the set $\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$ for any $x\in X$. For any $(\xi,v),(\zeta,w)\in\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$, we set
$d((\xi,v),(\zeta,w))=\inf_{\chi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})}\{d_{\chi}((\xi,v),(\zeta,w))\}+d_{X}(v,w).$
where $d_{\chi}((\xi,v),(\zeta,w)):=\max(d_{v}(\xi,\chi_{v}),d_{w}(\zeta,\chi_{w}))$. It can be checked that $d$ is a well-defined metric on $\operatorname{Hom}_{C^{*}}(A,B)$.
Proposition 3.10.
@@
$\bullet$ Application of the selection theorem. For the subsequent paragraph, we fix the following that we will refer to as the current setting.
We let $(X,\{A_{x}\}_{x},\mathcal{A})$ and $(X,\{B_{x}\}_{x},\mathcal{B})$ be continuous fields of separable $C^{*}$-algebras over the same compact Hausdorff metric space $(X,d_{X})$ of dimension at most one. We fix metrics $\{d_{x}\}_{x}$ and $d$ on $\{\operatorname{Hom}_{C^{*}}(A_{x},B_{x})\}_{x}$ and $\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x})$ respectively as in  A. (Recall that $d$ induces $\tau_{(\mathcal{A}\rightarrow\mathcal{B})}$ and $d_{x}$ induces $\tau_{pn}$.)
We let $\alpha\colon\operatorname{Cu}(\mathcal{A})\rightarrow\operatorname{Cu}(\mathcal{B})$ be any $\operatorname{Cu}$-morphism which descends to the fibers and such that $\alpha_{x}([1_{A_{x}}])=[1_{B_{x}}]$. We consider the multi-valued map
$\begin{array}[]{ll}\Theta_{\alpha}\colon X\longrightarrow\mathcal{S}_{\alpha}\\ \hskip 25.6073ptx\longmapsto\{\phi_{x}\in\operatorname{Hom}_{C^{*}}(A_{x},B_{x})\mid\operatorname{Cu}(\phi_{x})=\alpha_{x}\}\end{array}$
where $\mathcal{S}_{\alpha}:=\{\Theta_{\alpha}(x)\}_{x\in X}$ is seen as a family of subsets of $(\bigsqcup_{x}\operatorname{Hom}_{C^{*}}(A_{x},B_{x}),d)$.
Proposition 3.11.
Retain the current setting. Then the family $\mathcal{S}_{\alpha}$ satisfies condition (i) of Michael’s selection theorem. (See  A.)
Proof.
[…]
We now turn to the uniform path property of $\mathcal{S}_{\alpha}$ and the lower-semicontinuity of $\Theta_{\alpha}$. Before diving into the proof, we will need the following results.
3.12Admissible sets.
Recall that $\operatorname{Cu}$-morphisms $\alpha,\beta\colon S\rightarrow T$ compare on a (finite) set $G\in S$, and we write $\alpha\simeq_{G}\beta$, if $\alpha(g^{\prime})\leq\beta(g)$ and $\beta(g^{\prime})\leq\alpha(g)$ for any $g^{\prime},g\in G$ with $g^{\prime}\ll g$. (See [8, Definition 3.9].)
It has been shown in [22, Theorem 3.3.1], that the classification of *-homomorphisms between unital $\operatorname{AI}$-algebras (recalled in Section 3) is uniformly continuous. More specifically, given a unital $\operatorname{AI}$-algebra $A$, a finite set $F\subseteq A$ and some arbitrary precision $\epsilon>0$, there exists a finite set $G\subseteq\operatorname{Cu}(A)$ satisfying the following:
For any (unital) *-homomorphisms $\phi,\psi\colon A\rightarrow B$ into a unital $\operatorname{AI}$-algebra $B$, there exists $u\in U(B)$ such that $\phi\simeq_{(F,\epsilon)}u\psi u^{*}$, whenever $\operatorname{Cu}(\phi)\simeq_{G}\operatorname{Cu}(\psi)$.
We refer to $G$ as a $(F,\epsilon)$-admissible set for $A$.
@@
Retain the current setting. Assume there exists $x\in X$ with $\alpha_{x}=\beta_{x}$.
Then for any finite set $G\subseteq\operatorname{Cu}(\mathcal{A})$, there exists an open neighborhood $V$ of $x$ such that $\alpha_{v}\simeq_{G_{v}}\beta_{v}$ for all $v\in V$.
@@
By a trivial field, we mean a continuous field $(X,\{A_{x}\}_{x},\mathcal{A})$ such that $\mathcal{A}$ is isomorphic to $C(X)\otimes A$, for some $C^{*}$-algebra $A$.
By a locally trivial field, we mean a continuous field $(X,\{A_{x}\}_{x},\mathcal{A})$ satisfying that for any $x\in X$, there exists an open neighborhood $V$ of $x$ such that the (closed two-sided) ideal $\mathcal{I}_{U}:=C_{0}(U)\mathcal{A}$ is isomorphic to $C_{0}(U)\otimes A_{x}$, for some $C^{*}$-algebra $A_{x}$.
Theorem 3.14.
Retain the current setting. Assume moreover that $(X,\{A_{x}\}_{x},\mathcal{A})$ is a locally trivial field.
  • (i)
    The family $\mathcal{S}_{\alpha}$ satisfies the uniform path property. See  A (ii).
  • (ii)
    The map $\Theta_{\alpha}\colon X\longrightarrow\mathcal{S}_{\alpha}$ is lower-semicontinuous whenever $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ satisfies $\rm(PP)$.
Proof.
(i) Recall that $A_{x}$ is interval algebras for any $x\in X$. Combined with the (local) triviality of $(X,\{A_{x}\}_{x},\mathcal{A})$, we deduce the uniform path property from [6, Theorem 2.3] that we will explicitly recall and use later. (See  B).
(ii) Let $V_{\epsilon}^{\chi}$ be a basic open set of $\tau_{(\mathcal{A}\rightarrow\mathcal{B})}$. We aim to show that $\{x\in X\mid\Theta_{\alpha}(x)\cap V_{\epsilon}^{\chi}\neq\emptyset\}$ is open in $X$. Let $x\in X$ be such that $\Theta_{\alpha}(x)\cap V_{\epsilon}^{\chi}\neq\emptyset$. It suffices to prove that there exists an open neighborhood $Y$ of $x$ such that $Y\subseteq\{x\in X\mid\Theta_{\alpha}(x)\cap V_{\epsilon}^{\chi}\neq\emptyset\}$.
First, observe that $x\in V$ and that there exists $\phi_{x}\in\Theta_{\alpha}(x)$ such that $d_{x}(\phi_{x},\chi_{x})<\epsilon$. Choose $\epsilon^{\prime}>0$ such that $d_{x}(\phi_{x},\chi_{x})<\epsilon^{\prime}<\epsilon$. By $\rm(PP)$, we can find $\psi\in\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ such that $\psi_{x}=\phi_{x}$ and Lemma 3.6 yields an open neighborhood $W$ of $x$ such that $d_{w}(\psi_{w},\chi_{w})<\epsilon^{\prime}$ for all $w\in W$.
Now, fix $\delta>0$ small enough such that $\delta+\epsilon^{\prime}<\epsilon$ and a finite set $F\subseteq\mathcal{A}$. Let $x\in X$. By the local triviality of $(X,\{A_{x}\}_{x},\mathcal{A})$, we can find an open neighborhood $Z$ of $x$ such that $A_{z}\simeq A_{x}$ for all $z\in Z$. Now, we apply the uniform continuity of the classification recalled in  A, to find a finite set $G\subseteq\operatorname{Cu}(\mathcal{A})$, such that $G_{z}$ is a $(F_{z},\delta)$-admissible set for all $z\in Z$. Since $\operatorname{Cu}(\psi_{x})=\alpha_{x}$, we can apply Lemma 3.13 to find an open neighborhood $\tilde{Z}$ of $x$ such that $\operatorname{Cu}(\psi_{z})\simeq_{G_{z}}\alpha_{z}$ for all $z\in\tilde{Z}$. Also, recall that $\Theta_{\alpha}(z)$ is non-empty. Henceforth, for any $z\in Z\cap\tilde{Z}$, we are able to find $\phi_{z}\in\Theta_{\alpha}(z)$ such that $d_{z}(\psi_{z},\phi_{z})<\delta$.
Putting everything together, we compute that
$d_{z}(\phi_{z},\chi_{z})\leq d_{z}(\phi_{z},\psi_{z})+d_{z}(\psi_{z},\chi_{z})\leq\delta+\epsilon^{\prime}<\epsilon$
for any $z\in W\cap Z\cap\tilde{Z}$. As a conclusion, for any $y\in Y:=W\cap Z\cap\tilde{Z}\cap V$, we have $\Theta_{\alpha}(y)\cap V_{\epsilon}^{\chi}\neq\emptyset$ which ends the proof. ∎
As a consequence, we can invoke Michael’s selection theorem (see  A) to obtain the following existence result.
Corollary 3.15.
@@
A fortiori, $\operatorname{Cu}(\phi_{x})=\alpha_{x}$ for any $x\in X$.
@@
It is immediate that the triviality of $(X,\{A_{x}\}_{x},\mathcal{A})$ and $(X,\{B_{x}\}_{x},\mathcal{B})$ implies local triviality of $(X,\{A_{x}\}_{x},\mathcal{A})$ and $\rm(PP)$ of $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$. The result now follows from  A and  A. ∎
We remark that the restriction to trivial continuous fields was imposed solely to obtain $\rm(PP)$. Consequently, establishing $\rm(PP)$ for a class of locally trivial fields would lead to a more general existence result.
3.16.
We thank M. Dădărlat for pointing out that the Mayer-Vietoris sequence applied to $\mathrm{K}$-theory could provide an obstruction to $\rm(PP)$ for locally trivial fields over the circle. Let us construct a specific example.
@@
Finally, let $V,W$ be proper closed subintervals of $\mathbb{T}$ such that $V\cup W=\mathbb{T}$. It is well known that $\mathcal{A}\simeq\mathcal{A}(V)\oplus_{\mathcal{A}(V\cap W)}\mathcal{A}(W)$ and $\mathcal{B}\simeq\mathcal{B}(V)\oplus_{\mathcal{B}(V\cap W)}\mathcal{B}(W)$. (See e.g. [5] for the sheaves picture of continuous fields.) The Mayer-Vietoris sequence yields $\mathrm{K}_{0}(\mathcal{A})\simeq\mathrm{K}_{0}(\mathcal{B})$ which leads to a contradiction. As a consequence, $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ does not have $\rm(PP)$.
Question 3.17.
Let $\mathcal{A},\mathcal{B}$ be continuous fields of interval algebras over the interval. Under which conditions $\operatorname{Hom}_{\mathcal{F}_{C^{*}}}(\mathcal{A},\mathcal{B})$ has $\rm(PP)$?